# HU-EP-26/09

### v1

@@ 1 Introduction
The AdS soliton is a smooth, cigar-like geometry in the infrared obtained via a double Wick rotation of a black hole solution. This has been studied vastly in [12, 11, 32, 16, 14, 13, 41, 40, 31, 29, 30, 33, 9, 8, 10]. The soliton of [1] generalizes the supersymmetric soliton introduced in [2] by incorporating a non-trivial scalar profile, an additional gauge field, and an extra parameter. This wider class of solutions has been interpreted as describing Coulomb-branch deformations of $\mathcal{N}=4$ SYM [17, 22]. The corresponding five-dimensional gauged supergravity soliton and its uplift to type-IIB supergravity, further explored in [14, 13], is used as the seed for constructing new holographic duals via the TsT solution-generating technique. This seed background describes supersymmetry-preserving RG flows from a four-dimensional superconformal field theory (SCFT) in the UV to a three-dimensional quantum field theory in the IR. It exhibits confinement of external quarks and possesses a mass gap.
@@ 3 Observables
The marginal transformations only affect the internal part of the geometry. Consequently, some observables are expected to exhibit the same behaviour as in the case of the seed background [14, 13].
@@ 3.2 Wilson Loop
$\textbf{(A)}\quad\psi_{0}=\text{any const.},\,\theta_{0}=\frac{\pi}{2}\,,\quad\textbf{(B)}\quad\psi_{0}=\theta_{0}=0\,,\quad\textbf{(C)}\quad\psi_{0}=0\,,\theta_{0}=\frac{\pi}{2}\,.$ (3.16)
@@ 5 Acknowledgements
We are grateful to Carlos Nuñez for carefully reading the draft and providing useful comments and insights. We also thank Dimitrios Chatzis, Prem Kumar, Dimitrios Zoakos and Federico Castellani for discussions. M.H thanks Durham University YTF25 for hospitality and for the chance to speak about this work. M.H. has been supported by the STFC consolidated grant ST/Y509644/1. The research of G.I. is supported by the Einstein Stiftung Berlin via the Einstein International Postdoctoral Fellowship program “Generalised dualities and their holographic applications to condensed matter physics” (project number IPF- 2020-604).

### v2

@@ 1 Introduction
The AdS soliton is a smooth, cigar-like geometry in the infrared obtained via a double Wick rotation of a black hole solution. This has been studied vastly in [13, 12, 35, 17, 15, 14, 44, 43, 34, 31, 33, 36, 10, 9, 11, 49]. The soliton of [1] generalizes the supersymmetric soliton introduced in [2] by incorporating a non-trivial scalar profile, an additional gauge field, and an extra parameter. This wider class of solutions has been interpreted as describing Coulomb-branch deformations of $\mathcal{N}=4$ SYM [18, 23]. The corresponding five-dimensional gauged supergravity soliton and its uplift to type-IIB supergravity, further explored in [15, 14], is used as the seed for constructing new holographic duals via the TsT solution-generating technique [32]. This seed background describes supersymmetry-preserving RG flows from a four-dimensional superconformal field theory (SCFT) in the UV to a three-dimensional quantum field theory in the IR. It exhibits confinement of external quarks and possesses a mass gap.
@@ 3 Observables
In this section we study various observables for each of the TsT-deformed backgrounds of Sec. 2. These include the Page charges, Wilson loops, ’t Hooft loops, entanglement entropy, and the holographic central charge flow.
@@ 3.2 Wilson Loop
$\textbf{(A)}\quad\theta_{0}=\frac{\pi}{2}\,,\psi_{0}=\text{any const.},\quad\textbf{(B)}\quad\theta_{0}=\psi_{0}=0\,,\quad\textbf{(C)}\quad\theta_{0}=0\,,\psi_{0}=\frac{\pi}{2}\,.$ (3.16)
@@ 3.4 Entanglement Entropy
The presence of a phase transition in the entanglement entropy is not unique to confining geometries. As seen in [25], a phase transition in the entanglement entropy does not necessarily imply confinement. A TsT with the T-dual performed along the confining direction, $\phi$, will give a non-confining non-commutative geometry as the coefficient of $\mathrm{d}\phi^{2}$ will be deformed and the $S^{1}$ isometry is broken. However, this non-confining geometry will also have a phase transition in the entanglement entropy, it is the TsT which introduces a non-commutativity parameter [48],[32] which introduces a length parameter into the metric. This causes the holographic EE to saturate because of a finite correlation length in the system. This modifies the UV in the dual field theory and the entanglement entropy exhibits a cross-over from the area law to a volume law. See [5] for a similar study for non-commutative versions of the ABJM theory, where the TsT deformation generates UV modifications and a corresponding phase transition in the EE despite the absence of confinement.
@@ 5 Acknowledgements
We are grateful to Carlos Nuñez for carefully reading the draft and providing useful comments and insights. We also thank Dimitrios Chatzis, Prem Kumar, Dimitrios Zoakos, Niko Jokela and Federico Castellani for discussions. M.H thanks Durham University YTF25 for hospitality and for the chance to speak about this work. M.H. has been supported by the STFC consolidated grant ST/Y509644/1. The research of G.I. is supported by the Einstein Stiftung Berlin via the Einstein International Postdoctoral Fellowship program “Generalised dualities and their holographic applications to condensed matter physics” (project number IPF- 2020-604).